The Gentle Warming Company is purchasing a new cooling system. To pay for the system, the company agrees to make 12 installment payments of $6669 per year, beginning at the end of year 1. The company also has to pay a system fee of $2036 that will begin at the end of year 1 and decrease by $199 per year until the system is paid off. How much should the company set aside today to pay for the cooling system if the bank pays 3% per year compounded annually? Enter your answer as follows: 12345 Round your answer. Do not use a dollar sign ($), any commas () or a decimal point ().
Answer #1

Amount = installment+system fee

= 6669+2036

= 8705

P= 8705(P/A,3%,12) - 199(P/G,3%12)

= 8705(9.954) - 199(51.248)

= 86649.57 - 10198.352

= 76451